# Escaping Plato’s Cave: JAM for Aligning Independently Trained Vision and Language Models

### v5

@@ 1.3 Related Work
Testing for Alignment. Prior work in this direction has been largely diagnostic focusing on evaluating alignment between frozen features with broad, context-agnostic datasets (e.g., Wikipedia caption dataset (WIT) [4]): measures such as centered kernel alignment (CKA) [5], variants of CCA (SVCCA [6], projection-weighted CCA [7]), and nearest-neighbor metrics [8, 1]. Other approaches explore probing tasks or zero-shot transfer to assess latent compatibility across modalities [9, 10]. More recently, model-stitching frameworks have examined whether independently trained sub-networks can be functionally composed to perform new tasks [3], hinting at deeper interoperability between pretrained systems but focusing on vision domain [3, 11].
@@ 3 Our method: Joint Autoencoder Modulator (JAM)
Our approach learns a joint autoencoder across two disjoint pre-trained models, as shown in Figure 3. We first utilize a modality-specific autoencoder for each pre-trained model which distills $d_{input}\rightarrow d_{latent}$, which can be viewed as a form of statistical regularisation reminiscent of PCA [25, 26]: the network is encouraged to find a compact set that preserve essential variance while discarding noise through reconstruction loss (i.e., MSE Loss). The joint training objective then contrastively pulls the latent vectors from matched image–text pairs together and push hard-negatives apart. Thus, the respective autoencoders have two objectives: denoising for each modality as well as optimizing for cross-modal relation (the exact loss functions below).
@@
The encoder follows a “funnel” layout: a sequence of fully connected layers that progressively reduce the dimensionality of the input embedding, each stage wrapped in LayerNorm for stable statistics [27], a SwiGLU non-linearity for gated expressiveness [28], and dropout for regularisation. After every dense layer, a lightweight MLP residual block [29] re-injects the intermediate representation back into itself, ensuring gradient flow. Passing through these hidden layers, in a bottleneck stage, a final linear projection maps the hidden width to the fixed latent size. The decoder mirrors the encoder, walking back through the hidden sizes in reverse. Together, we chose this design to be an effective framework for building compact yet powerful module that fits both high-dimensional text and vision representations (i.e., embeddings). In our experiments, we used 3 hidden layers with dimension size of 512, and bottleneck dimension size as 256.
@@ 3.1 Loss functions
$\begin{split}\mathcal{L}_{\text{VL}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\sigma(v_{i},l_{P_{i}})}{\sum_{j=1,j\neq i}^{N}\sigma(v_{i},l_{P_{j}})},\quad\mathcal{L}_{\text{LV}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\sigma(l_{P_{i}},v_{i})}{\sum_{j=1,j\neq i}^{N}\sigma(l_{P_{i}},v_{j})}\end{split}$ (1)
@@ Appendix A Impact Statement
We propose JAM, a post-hoc adaptor for platonic alignment that:
  • Preserves specialization – Modality-specific autoencoders reconstruct each latent space, guarding task-relevant features.
  • Enforces cross-modal coherence – Our novel spread loss couples the JAM framework, amplifying the latent commonality predicted by the Platonic hypothesis.
  • Light-weight framework – no end-to-end multimodal fine-tuning (nor pre-training); refer to D.2 that shows the lightweight structure of JAM.
@@
  • Platonic Alignment through Specialist Task Settings: We test the platonic alignment through both statistical and model-training framework in specialist task settings which require more fine-grained contextual understanding of the data.
@@ B.2 Metrics Formulation
Let $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$ be paired samples: in our case, let $x_{i}$ be the images and $y_{i}$ be the corresponding texts. A representation function $f_{V}:\mathcal{X}\!\to\!V\in\mathbb{R}^{d}$ maps inputs to feature vectors $v_{i}=f_{V}(\mathbf{x}_{i})$ in an RKHS (Reproducing Kernel Hilbert Space) $\mathcal{H}$ [52] equipped with the inner-product kernel $K_{ij}=\kappa(v_{i},v_{j})=\langle v_{i},v_{j}\rangle$. Analogously $f_{L}$ maps $\mathbf{y}_{i}$ to $l_{i}$ with kernel $L_{ij}=\langle l_{i},l_{j}\rangle$. Center both kernels to remove mean effects via the centring matrix $H=I-\frac{1}{n}\mathbf{1}\mathbf{1}^{\!\top}$:
@@
B.2.1 Centered-Kernel Alignment (CKA)
@@
Because the denominator applies the same operation to each modality, CKA is scale-invariant and bounded in $[0,1]$; it therefore has advantage in measuring global geometric correspondence. However, equation (5) aggregates all pairwise relations, so even a small region of mis-aligned samples suppresses the score which CKA can be too strict for measuring cross-modal cases.
B.2.2 Centered-Kernel $k$-NN Alignment (CKNNA)
To focus on local structure, the authors in [8, 1] suggest a binary mask $\alpha(i,j)$ that selects only pairs that are mutual $k$-nearest neighbours in both modalities:
@@ B.2.2 Centered-Kernel $k$-NN Alignment (CKNNA)
Using this mask, one can restrict the definition of alignment to bias more towards the local structure through replacing the full HSIC trace by a masked cross-covariance.:
@@ B.2.3 CCA (Canonical Correlation Analysis)
Let $X\!\in\!\mathbb{R}^{d_{x}\times n}$ and $Y\!\in\!\mathbb{R}^{d_{y}\times n}$ be column-wise zero-centered feature matrices produced by two representations (all means are removed so that covariance estimates are unbiased). Define sample covariance blocks:
@@
with $P_{x}\!\in\!\mathbb{R}^{d_{x}\times r}$ (orthonormal) obtained from linear PCA or kernel PCA using an RBF kernel $\kappa(\mathbf{z},\mathbf{z}^{\prime})=\exp[-\gamma\|\mathbf{z}-\mathbf{z}^{\prime}\|^{2}]$. CCA is then run on $\tilde{X},\tilde{Y}$, producing more stable correlations in high-dimensional regimes. In our experiments, we used $r=50$ (i.e., feature dimension = 50), and $k=50$ (i.e., number of canonical components = 50).
@@ B.2.4 SVCCA (Singular Vector Canonical Correlation Analysis)
SVCCA [6] replaces the heuristic PCA cut-off with a data-dependent approach via SVD:
@@ C.2 Spread Loss Formulation in Visual Representation Learning Domain
C.3 Side-by-Side Comparison
@@
$\mathcal{L}_{ConCon}$ pulls an image toward all texts that share its visual context (analogous to class-level clustering, performed by $\mathcal{L}_{sup}$), while $\mathcal{L}_{contextNCE}$ pushes it away from other texts inside that context except its primary caption (analogous to instance discrimination, performed by $\mathcal{L}_{cNCE}$).
Table 10 summarizes the symbols used in the Spread loss in vision domain (left column) and their direct counterparts in our our multimodal formulation (right column). The first row clarifies that the visual setting relies on a single encoder $f(\cdot)$, whereas the multimodal variant distinguishes between an image encoder $\Phi_{V}$ and a text encoder $\Phi_{L}$. Subsequent rows pair up the anchor, positive, and negative sets, making it explicit that class labels in the visual domain translate to visual contexts (collections of captions). Finally, the similarity function retains the same softmax-temperature structure; only the argument types differ (image-image vs. image–text).
@@
$\mathcal{L}_{sup}$ $x_{i}$ same class other classes intra-class cohesion
$\mathcal{L}_{ConCon}$ $v_{i}$ $C_{L}(i)$ $\widetilde{C}_{L}(i)$ cross-modal cohesion

### v6

@@ 1.3 Related Work
Testing for Alignment. Prior work in this direction has been largely diagnostic focusing on evaluating alignment between frozen features with broad, context-agnostic datasets (e.g., Wikipedia caption dataset (WIT) [45]): measures such as centered kernel alignment (CKA) [29], variants of CCA (SVCCA [42], projection-weighted CCA [36]), and nearest-neighbor metrics [28, 22]. Other approaches explore probing tasks or zero-shot transfer to assess latent compatibility across modalities. More recently, model-stitching frameworks have examined whether independently trained sub-networks can be functionally composed to perform new tasks [4], hinting at deeper interoperability between pretrained systems but focusing on vision domain [4, 30].
@@ 3 Our method: Joint Autoencoder Modulator (JAM)
Our approach learns a joint autoencoder across two disjoint pre-trained models, as shown in Figure 3. We first utilize a modality‑specific autoencoder for each pre-trained model which distills $d_{input}\rightarrow d_{latent}$, which can be viewed as a form of statistical regularisation reminiscent of PCA [2, 5]: the network is encouraged to find a compact set that preserve essential variance while discarding noise through reconstruction loss (i.e., MSE Loss). The joint training objective then contrastively pulls the latent vectors from matched image–text pairs together and push hard‑negatives apart. Thus, the respective autoencoders have two objectives: denoising for each modality as well as optimizing for cross-modal relation (the exact loss functions below).
@@
The encoder follows a “funnel” layout: a sequence of fully connected layers that progressively reduce the dimensionality of the input embedding, each stage wrapped in LayerNorm for stable statistics [3], a SwiGLU non‑linearity for gated expressiveness [44], and dropout for regularisation. After every dense layer, a lightweight MLP residual block [49] re‑injects the intermediate representation back into itself, ensuring gradient flow. Passing through these hidden layers, in a bottleneck stage, a final linear projection maps the hidden width to the fixed latent size. The decoder mirrors the encoder, walking back through the hidden sizes in reverse. Together, we chose this design to be an effective framework for building compact yet powerful module that fits both high‑dimensional text and vision representations (i.e., embeddings). In our experiments, we used 3 hidden layers with dimension size of 512, and bottleneck dimension size as 256.
@@ 3.1 Loss functions
$\begin{split}\mathcal{L}_{\text{VL}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\sigma(v_{i},l_{P_{i}})}{\sum_{j=1,j\neq i}^{N}\sigma(v_{i},l_{P_{j}})},\qquad\mathcal{L}_{\text{LV}}=-\frac{1}{N}\sum_{i=1}^{N}\log\frac{\sigma(l_{P_{i}},v_{i})}{\sum_{j=1,j\neq i}^{N}\sigma(l_{P_{i}},v_{j})}\end{split}$ (1)
@@ Appendix A Impact Statement
We propose JAM, a post‑hoc adaptor for platonic alignment that:
  • Preserves specialization – Modality‑specific autoencoders reconstruct each latent space, guarding task‑relevant features.
  • Enforces cross‑modal coherence – Our novel spread loss couples the JAM framework, amplifying the latent commonality predicted by the Platonic hypothesis.
  • Light-weight framework – no end‑to‑end multimodal fine‑tuning (nor pre-training); refer to D.2 that shows the lightweight structure of JAM.
@@
  • Platonic Alignment through Specialist Task Settings: We test the platonic alignment through both statistical and model-training framework in specialist task settings which require more fine‑grained contextual understanding of the data.
@@ B.2 Metrics Formulation
Let $\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{n}$ be paired samples: in our case, let $x_{i}$ be the images and $y_{i}$ be the corresponding texts. A representation function $f_{V}:\mathcal{X}\!\to\!V\in\mathbb{R}^{d}$ maps inputs to feature vectors $v_{i}=f_{V}(\mathbf{x}_{i})$ in an RKHS (Reproducing Kernel Hilbert Space) $\mathcal{H}$ [14] equipped with the inner‑product kernel $K_{ij}=\kappa(v_{i},v_{j})=\langle v_{i},v_{j}\rangle$. Analogously $f_{L}$ maps $\mathbf{y}_{i}$ to $l_{i}$ with kernel $L_{ij}=\langle l_{i},l_{j}\rangle$. Center both kernels to remove mean effects via the centring matrix $H=I-\frac{1}{n}\mathbf{1}\mathbf{1}^{\!\top}$:
@@
B.2.1 Centered‑Kernel Alignment (CKA)
@@
Because the denominator applies the same operation to each modality, CKA is scale‑invariant and bounded in $[0,1]$; it therefore has advantage in measuring global geometric correspondence. However, equation (5) aggregates all pairwise relations, so even a small region of mis‑aligned samples suppresses the score which CKA can be too strict for measuring cross-modal cases.
B.2.2 Centered‑Kernel $k$-NN Alignment (CKNNA)
To focus on local structure, the authors in [28, 22] suggest a binary mask $\alpha(i,j)$ that selects only pairs that are mutual kk‑nearest neighbours in both modalities:
@@ B.2.2 Centered‑Kernel $k$-NN Alignment (CKNNA)
Using this mask, one can restrict the definition of alignment to bias more towards the local structure through replacing the full HSIC trace by a masked cross‑covariance.:
@@ B.2.3 CCA (Canonical Correlation Analysis)
Let $X\!\in\!\mathbb{R}^{d_{x}\times n}$ and $Y\!\in\!\mathbb{R}^{d_{y}\times n}$ be column‑wise zero-centered feature matrices produced by two representations (all means are removed so that covariance estimates are unbiased). Define sample covariance blocks:
@@
with $P_{x}\!\in\!\mathbb{R}^{d_{x}\times r}$ (orthonormal) obtained from linear PCA or kernel PCA using an RBF kernel $\kappa(\mathbf{z},\mathbf{z}^{\prime})=\exp[-\gamma\|\mathbf{z}-\mathbf{z}^{\prime}\|^{2}]$. CCA is then run on $\tilde{X},\tilde{Y}$, producing more stable correlations in high‑dimensional regimes. In our experiments, we used $r=50$ (i.e., feature dimension = 50), and $k=50$ (i.e., number of canonical components = 50).
@@ B.2.4 SVCCA (Singular Vector Canonical Correlation Analysis)
SVCCA [42] replaces the heuristic PCA cut‑off with a data‑dependent approach via SVD:
@@ C.2 Spread Loss Formulation in Visual Representation Learning Domain
C.3 Side‑by‑Side Comparison
@@
$\mathcal{L}_{ConCon}$ pulls an image toward all texts that share its visual context (analogous to class‑level clustering, performed by $\mathcal{L}_{sup}$), while $\mathcal{L}_{contextNCE}$ pushes it away from other texts inside that context except its primary caption (analogous to instance discrimination, performed by $\mathcal{L}_{cNCE}$).
Table 10 summarizes the symbols used in the Spread loss in vision domain (left column) and their direct counterparts in our our multimodal formulation (right column). The first row clarifies that the visual setting relies on a single encoder $f(\cdot)$, whereas the multimodal variant distinguishes between an image encoder $\Phi_{V}$ and a text encoder $\Phi_{L}$. Subsequent rows pair up the anchor, positive, and negative sets, making it explicit that class labels in the visual domain translate to visual contexts (collections of captions). Finally, the similarity function retains the same softmax‐temperature structure; only the argument types differ (image-image vs. image–text).
@@
$\mathcal{L}_{sup}$ $x_{i}$ same class other classes intra‑class cohesion
$\mathcal{L}_{ConCon}$ $v_{i}$ $C_{L}(i)$ $\widetilde{C}_{L}(i)$ cross‑modal cohesion